OpenStudy (anonymous):

0! = 0 Is this true or false? And what does the "!" mean?

OpenStudy (highschoolmom2010):

The statement A! is true if and only if A is false

OpenStudy (anonymous):

?

OpenStudy (anonymous):

"!" is the symbol for a factorial factorials are defined like this \[n! = n * (n-1) * (n-2) * ... *1\] where n is a positive number. you can think of factorials as the number of permutations of numbers from 1 to n, EX 2! = 2 * 1 = 2 so how many ways can you write {1, 2} ? {1, 2} , {2, 1} , so 2 ways 1! = 1 {1}, 1 way 0! is tricky, you need to get into sets, but there is a thing called a null set, a set that contains nothing {}, and you can write this 1 way therefore, 0! = 1

OpenStudy (anonymous):

another example of a factorial, 6! = 6 * 5 * 4 * 3 * 2 * 1, 7! = 7 * 6 * 5 * 4 * 3 * 2 * 1

OpenStudy (anonymous):

Ohh ok, thanks :)